# VISTA: Knowledge-Driven Interpretable Vessel Trajectory Imputation via Large Language Models

### v1

@@ 3.3.1. Incomplete AIS Data
Example 0.
Continuing Example 3.7, using the convention $M_{k}=0$ for a segment with missing records and $M_{k}=1$ for a fully observed segment, the configuration in Figure 4 yields $M_{k-1}=1$, $M_{k}=0$, and $M_{k+1}=0$, i.e., the local portion of the mask vector $\mathbf{M}_{\iota}$ is $\langle\ldots,1,0,0,\ldots\rangle$.
3.3.2. Updating of Structured Data-derived Knowledge
A minimal segment $\mathbf{S}_{\iota}^{k}$ is a time-indexed sample and cannot directly update the SD-KG $\mathcal{G}_{d}$, which is organized over semantic node types and their relations. To bridge segment-level data and the graph-level representation, we introduce the knowledge unit as the minimal semantic unit distilled from a segment: it aligns static attribute, behavior pattern, and imputation function, serving as the atomic granularity for SD-KG updating.
Definition 0 (Knowledge Unit).
A knowledge unit $u$ is defined as a triple $u\;=(v_{s},v_{b},v_{f})\;=((\iota,\eta,\chi,d,\ell,\beta,\kappa,\sigma),\;(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau}),\;\\ (f,d(f)))$, where $v_{s}$ is static attributes, $v_{b}$ is a behavior pattern, and $v_{f}$ is an imputation method.
Given a minimal segment $\mathbf{S}_{\iota}^{k}$ with $M_{\iota}^{k}=1$, the updating of SD-KG refers to the procedure
  • Knowledge distillation: $\mathbf{S}_{\iota}^{k}\mapsto u_{\iota}^{k}=(v_{s},v_{b},v_{f})$, which compresses the raw segment into static attributes $v_{s}$, behavior pattern $v_{b}$, and imputation method $v_{f}$.
  • Graph integration and maintenance: The node sets $\mathcal{V}_{d}$ and edge sets $\mathcal{E}_{d}$ of $\mathcal{G}_{d}$ are updated by inserting or merging $v_{s},v_{b},v_{f}$ and adjusting the corresponding edges $(v_{s},v_{b})\in\mathcal{E}_{sb}$ and $(v_{b},v_{f})\in\mathcal{E}_{bf}$. Redundant nodes are merged and edge weights aggregated to ensure that $\mathcal{G}_{d}$ remains compact and efficient for retrieval.
Example 0.
As illustrated in Figure 5, there is a segment $\mathbf{S}_{\iota}^{k}=\langle x_{t},\ldots,x_{t+m-1}\rangle$ with $M_{\iota}^{k}=1$. Within this segment, the speed $s$ remains stable at approximately $5.4$ kn, the course $\theta$ varies gradually around $110^{\circ}$, the heading $\psi$ exhibits mild fluctuations near $110^{\circ}$, and both the navigation status and vessel type are constant, recorded as “under way using engine” and “cargo”.
Step 1: Knowledge distillation. From the static and status-related attributes, together with the vessel’s geographic position, we derive the static attributes $v_{s}=(\iota,\eta,\kappa,\sigma)$, describing a cargo vessel $\iota$ operating “under way using engine” near the Traffic Separation Scheme in the Great Belt East Channel. From the kinematic attributes, we extract the behavior pattern $v_{b}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$, corresponding to stable speed, gradual course change, mild heading fluctuation, lane-following intent, and duration 1300130013501350s. To avoid redundancy in $\mathcal{V}_{b}$, the states $p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau}$ are preferentially aligned with existing entries, in this case matching $v_{b,3}\in\mathcal{V}_{b}$ (continuing Example 3.5). Finally, based on the trajectory trend, we generate the imputation method $v_{f}=(f,d(f))$, where $f$ is the execute function that simulates the observed dynamics and $d(f)$ records its description.
Step 2: Graph integration and maintenance. Continuing Example 3.5, $v_{b}$ coincides with $v_{b,3}$ and the imputation function $v_{f}$ is semantically equivalent to $v_{f,1}$. The graph is updated accordingly: the weights of edges $(v_{\eta,3},v_{b,3})$ and $(v_{\kappa,2},v_{b,3})$ are each incremented by $1$ (from $6$ to $7$, red edges), while a new edge $(v_{b,3},v_{f,1})$ is added with weight $1$ (blue edge).
Further details on the updating mechanism of SD-KG are presented in Section 4.2.
3.3.3. Knowledge-Driven Interpretable Trajectory Imputation
Before specifying the imputation procedure, we first formalize the target output. The outcome should not only reconstruct the missing trajectory but also incorporate supporting knowledge that justifies the decision process. This leads to the following definition.
Definition 0 (Knowledge-supported Imputation Outcome).
For a minimal segment $\mathbf{S}_{\iota}^{k}$ with $M_{\iota}^{k}=0$, a knowledge-supported imputation outcome is defined as $R_{\iota}^{k}=(\widehat{\mathbf{S}}_{\iota}^{k},\;\mathcal{J}_{\iota}^{k})$ where $\widehat{\mathbf{S}}_{\iota}^{k}$ denotes the imputed trajectory segment, and $\mathcal{J}_{\iota}^{k}$ represents the accompanying justification. The justification $\mathcal{J}_{\iota}^{k}$ consists of:
  • Behavior-pattern rationale $(v_{b}^{*},\,\mathcal{J}_{\iota}^{k,b})$: $v_{b}^{*}$ is the selected behavior pattern for $\mathbf{S}_{\iota}^{k}$, and $\mathcal{J}_{\iota}^{k,b}$ provides contextual and statistical evidence from $\mathcal{G}_{d}$ supporting this selection.
  • Function-selection rationale $(v_{f}^{*},\,\mathcal{J}_{\iota}^{k,f})$: $v_{f}^{*}$ is the selected imputation function applied to obtain $\widehat{\mathbf{S}}_{\iota}^{k}$, and $\mathcal{J}_{\iota}^{k,f}$ contains statistical evidence from $\mathcal{G}_{d}$ supporting the choice of this function.
  • Human-friendly explanation $\mathcal{J}_{\iota}^{k,h}$: a natural-language interpretation, which provides high-level reasoning on why the vessel may exhibit the chosen behavior.
Example 0.
Continuing Example 1.1, the imputed segment is $\widehat{\mathbf{S}}_{\iota}^{k}$, reconstructed as a smooth curve bridging the missing interval, as shown in upper-left in Figure 1. The selected behavior pattern $v_{b}^{*}$ is a decelerate-then-align maneuver with low heading fluctuation and queue-following intent, while its supporting evidence $\mathcal{J}_{\iota}^{k,b}$ indicates that this pattern occurs in about $68\%$ of comparable gaps in historical AIS data. The chosen imputation function $v_{f}^{*}$ is “CTR+SD”, and $\mathcal{J}_{\iota}^{k,f}$ records statistical support that over $70\%$ of such gaps are best reconstructed by this function. Finally, $\mathcal{J}_{\iota}^{k,h}$ provides a natural-language explanation: the maneuver arises from port-entry procedures, where vessels reduce speed and adjust heading to merge into inbound lanes under traffic separation rules.
@@ 3.3.3. Knowledge-Driven Interpretable Trajectory Imputation
Given a minimal segment $\mathbf{S}_{\iota}^{k}$ with $M_{\iota}^{k}=0$ and contextual knowledge units $(u_{\iota}^{-},u_{\iota}^{+})$, the knowledge-driven trajectory imputation produces a knowledge-supported imputation outcome
(8) $R_{\iota}^{k}\;\leftarrow\;\mathrm{Impute}(\mathbf{S}_{\iota}^{k},(u_{\iota}^{-},u_{\iota}^{+}),\mathcal{G}_{d},\mathcal{K}_{\ell}),$
where $u_{\iota}^{-}$ and $u_{\iota}^{+}$ denote the contextual knowledge units extracted from the nearest complete segments preceding and succeeding $\mathbf{S}_{\iota}^{k}$, respectively. The procedure consists of:
@@
  • Human-friendly explanation: Given the selected behavior pattern and imputation function $(v_{b}^{*},v_{f}^{*})$, along with the induced subgraph $\mathcal{G}_{d}[\mathcal{V}_{s}(\iota)\cup\{v_{b}^{*}\}\cup\{v_{f}^{*}\}]$ of $\mathcal{G}_{d}$ (where $\mathcal{V}_{s}(\iota)$ denotes vessel-specific statistics), an explanatory generation step produces the natural-language justification $\mathcal{J}_{\iota}^{k,h}$, providing high-level reasoning for the reconstructed behavior.
Further details on knowledge driven trajectory imputation are provided in Section 4.3.
@@
4. Knowledge-Driven Interpretable Vessel Trajectory Imputation via LLMs
@@ 4.2. SD-KG Construction and Maintenance
We construct and maintain the SD-KG through three modules followed by a graph update. For each complete minimal segment, the Static and Spatial Encoder produces static attributes and spatial context, the Behavior Abstraction maps motion signals to a behavior pattern, and the Imputation Method Builder registers or generates a validated executable method with a concise description. The outputs form a knowledge unit $u=\langle v_{s},v_{b},v_{f}\rangle$ as defined in Definition 3.10. Each knowledge unit is then integrated into the SD-KG through node alignment, and edge updates.
@@ 4.2.1. Static and Spatial Encoder
The encoder organizes the static attribute $v_{s}=\langle\iota,\eta,\chi,d,\ell,\beta,\sigma,\kappa\rangle$ into three classes and applies targeted processing. Discrete fields $\{\iota,\eta,\chi,\kappa\}$ are standardized, continuous fields $\{d,\ell,\beta\}$ are discretized into intervals, and spatial context $\{\sigma\}$ is inferred from geospatial surroundings. This design keeps the $\mathcal{V}_{s}$ in SD-KG compact, interpretable, and easy to maintain.
@@
Continuous Attributes. Draught $d$, length $\ell$, and width $\beta$ are discretized into interval-based categories to form finite vocabularies $\mathcal{V}_{d},\mathcal{V}_{\ell},\mathcal{V}_{\beta}$ and preserve interpretability. Specifically, $d$ is discretized into 2 m intervals up to 12 m, $\ell$ into 50 m intervals up to 300 m, and $\beta$ into 5 m intervals up to 30 m, each with an open-ended bin beyond the upper bound. These cutoffs reflect common maritime conventions (Authority, 2022; Clarksons, 2025) (e.g., Panamax thresholds at $\ell\approx 300$ m, $\beta\approx 32$ m, and $d\approx 12$ m) while maintaining sufficient support.
@@ 4.2.2. Behavior Abstraction
This component converts motion signals within the minimal segment into a discrete, human interpretable behavior pattern $v_{b}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$ and constrains vocabulary growth for interpretability and efficient retrieval.
Kinematic Tokens $(p^{s},p^{\theta},p^{\psi})$. Speed $s$, course over ground $\theta$, and heading $\psi$ exhibit diverse dynamics; direct tokenization yields opaque categories and uncontrolled growth. An LLM assisted procedure analyzes the temporal evolution in $\mathbf{S}_{\iota}^{k}$ and maps them to discrete variables $(p^{s},p^{\theta},p^{\psi})$. Vocabulary size is controlled by finite sets $\mathcal{P}^{s},\mathcal{P}^{\theta},\mathcal{P}^{\psi}$ with a deduplication rule that reuses semantically equivalent tokens before adding new ones.
@@
Duration Token ($p^{\tau}$). Duration token $p^{\tau}$ is computed from the segment time span and discretized into fixed width 50 s intervals, with an open ended bin for long gaps (i.e., $3600,+\infty)$).
@@ 4.2.3. Imputation Method Builder
The Imputation Method Builder produces the imputation method $v_{f}=(f,d(f))$, where $f$ is executable function and $d(f)$ is a concise usage description. Given a minimal segment $\mathbf{S}_{\iota}^{k}$, $v_{f}$ is generated through a three-stage process.
Function Proposal. Based on $\mathbf{S}_{\iota}^{k}$ and the associated $(v_{s},v_{b})$, we first retrieve the most relevant imputation method $v_{f}\in\mathcal{V}_{f}$ using the retrieval scheme described in Section 4.3.2, to prevent unnecessary duplication and keep $\mathcal{V}_{f}$ compact. If no suitable candidate is found, a LLM-based function generator produces a new executable function $f$ that can be directly applied to fit the observed trajectory within the segment.
Execution and Validation. The proposed function $f$ is executed on $\mathbf{S}_{\iota}^{k}$. At each timestamp we compute the absolute errors in latitude and longitude, then average them to obtain the overall fitting error $e(f)=\tfrac{1}{2}(\mathrm{MAE}_{\phi}+\mathrm{MAE}_{\lambda})$. The function is accepted if $e(f)$ does not exceed a predefined threshold of $\varrho_{f}=3e-3$. If the error exceeds this threshold, LLM-based function generator is provided with diagnostic feedback (error summaries), and asked to refine its proposal. This is repeated at most $\varrho_{r}$ times (default 3).
@@ 4.2.4. SD-KG Integration and Maintenance
Node Integration. Since static attributes, behavior patterns, and imputation methods have already undergone de-redundancy during knowledge unit extraction, the integration process is straightforward. For each static attribute, if a newly generated token does not yet exist, a corresponding node is created and added to $\mathcal{V}_{s}$. Similarly, if the behavior pattern $v_{b}$ and $v_{f}$ is not present in $\mathcal{V}_{b}$ and $\mathcal{V}_{f}$, a new node is created and inserted.
Edge Maintenance. For each knowledge unit $u=\langle v_{s},v_{b},v_{f}\rangle$, edges $(v_{s},v_{b})$ and $(v_{b},v_{f})$ are updated. If the edge does not exist, it is created with an initial weight of $1$. If it already exists, the weight is incremented by $1$, capturing the frequency of observed co-occurrences. This ensures that the SD-KG preserves both structural completeness and statistically grounded associations while remaining compact and retrievable.
@@ 4.3. Knowledge-Driven Trajectory Imputation
With the SD-KG $\mathcal{G}_{d}$ in place, we formalize knowledge driven trajectory imputation for an incomplete minimal segment $\mathbf{S}_{\iota}^{k}$. The procedure comprises three components: Behavior Estimator, which uses contextual constraints to query $\mathcal{G}_{d}$, rank candidates, and select the behavior pattern with supporting evidence; Method Selector, which retrieves an imputation method aligned with the chosen pattern, and executes it to reconstruct the gap; and Explanation Composer, which derives a human friendly explanation by combining a regulatory rule cue with an operational protocol rationale from the induced subgraph of SD-KG and contextual informations. In what follows, we detail each component.
@@ 4.3.1. Behavior Estimator
It operates in two steps: Candidate Set Generation, which queries the SD-KG to produce a ranked shortlist of plausible behaviors, and LLM-driven Selection and Rationale, which chooses the final pattern and explains the choice.
Candidate Set Generation. Given an incomplete segment $\mathbf{S}_{\iota}^{k}$ with context $(u_{\iota}^{-},u_{\iota}^{+})$, it infers the vessel-specific static attribute set $\mathcal{V}^{k}_{s}(\iota)$ from $(u_{\iota}^{-},u_{\iota}^{+})$, due to static attributes vary slowly over time. It then queries the SD-KG to retrieve behavior candidates via static–behavior edges
@@
and assign each candidate a graph prior based on normalized multiplicative support
(10) $\pi(v_{b})\;=\;\frac{\displaystyle\prod_{v\in\mathcal{V}^{k}_{s}(\iota)}\!(w_{sb}(v,v_{b})+1)}{\displaystyle\sum_{v_{b}^{\prime}\in\mathcal{C}_{b}}\ \prod_{v\in\mathcal{V}^{k}_{s}(\iota)}\!(w_{sb}(v,v_{b}^{\prime})+1)},$
where $w_{sb}$ is the weight in edge $(v,v_{b},w)$. It keeps the top-$K$ candidates by $\pi(v_{b})$ as the shortlist $\mathcal{C}_{b}^{K}$ (default K=5K{=}5).
LLM-driven Selection and Rationale. It first constructs the induced subgraph $\mathcal{G}_{\iota,b}^{K}=\mathcal{G}_{d}[\mathcal{V}^{k}_{s}(\iota)\cup\mathcal{C}_{b}^{K}]$, serializes it in a graph description language (i.e., DOT (Wikipedia contributors, 2025a)), and supplies it together with the contextual behavior patterns $v_{b}^{-}$ and $v_{b}^{+}$ extracted from $u_{\iota}^{-}$ and $u_{\iota}^{+}$. It then selects a final pattern $v_{b}^{*}\in\mathcal{C}_{b}^{K}$ based on the LLM’s analysis of these information.
Alongside the selection, it also obtains a two-part rationale: i) graph support, which highlights the most informative edges in $\mathcal{G}_{\iota,b}^{K}$ and their weights that favor the choice; and ii) contextual justification, which explains the consistency with $v_{b}^{-}$, $v_{b}^{+}$, and the gap’s boundary conditions. These two items constitute $\mathcal{J}_{\iota}^{k,b}$.

### v2

@@ 3.3.3. Knowledge-Driven Trajectory Imputation
Given a minimal segment $\mathbf{S}_{\iota}^{k}$ with $M_{\iota}^{k}=0$ and contextual knowledge units $(u_{\iota}^{-},u_{\iota}^{+})$, knowledge-driven trajectory imputation produces an imputation outcome with repair provenance $R_{i}^{k}$:
(7) $R_{\iota}^{k}\;\leftarrow\;\mathrm{Impute}(\mathbf{S}_{\iota}^{k},(u_{\iota}^{-},u_{\iota}^{+}),\mathcal{G}_{d}),$
where $u_{\iota}^{-}$ and $u_{\iota}^{+}$ are the contextual knowledge units extracted from the nearest complete segments preceding and succeeding $\mathbf{S}_{\iota}^{k}$, respectively. The procedure consists of:
@@
  • Causal explanation: Given the selected behavior pattern and imputation function $(v_{b}^{*},v_{f}^{*})$, along with the induced subgraph $\mathcal{G}_{d}[\mathcal{V}_{s}(\iota)\cup\{v_{b}^{*}\}\cup\{v_{f}^{*}\}]$ (where $\mathcal{V}_{s}(\iota)$ denotes the static attribute nodes of vessel $\iota$), an LLM generates the natural-language explanation $\mathcal{J}_{\iota}^{k,h}$, providing causal reasoning for the reconstructed behavior grounded in SDK context.
Further details on knowledge-driven trajectory imputation are provided in Section 4.3.
@@
4. Knowledge-Driven Interpretable Vessel Trajectory Imputation
@@ 4.2. SD-KG Construction and Maintenance
We construct and maintain the SD-KG through three components followed by a graph update. For each complete minimal segment, the Static and Spatial Encoder produces static attributes and spatial context, the Behavior Abstraction maps motion signals to a behavior pattern, and the Imputation Method Builder registers or generates a validated executable method with a concise description. The outputs form a knowledge unit $u=(v_{s},v_{b},v_{f})$ as defined in Definition 3.9. Each knowledge unit is then integrated into the SD-KG through node alignment and edge-weight updates.
@@ 4.2.1. Static and Spatial Encoder
The encoder organizes the static attribute $v_{s}=(\iota,\eta,\chi,d,\ell,\beta,\kappa,\sigma)$ into three classes and applies targeted processing. Discrete fields $\{\iota,\eta,\chi,\kappa\}$ are standardized, continuous fields $\{d,\ell,\beta\}$ are discretized into intervals, and spatial context $\{\sigma\}$ is inferred from geospatial surroundings. This design keeps the $\mathcal{V}_{s}$ in SD-KG compact, interpretable, and easy to maintain.
@@
Continuous Attributes. Draught $d$, length $\ell$, and width $\beta$ are discretized into interval-based categories to form finite vocabularies $\mathcal{V}_{d}$, $\mathcal{V}_{\ell}$, and $\mathcal{V}_{\beta}$ and preserve interpretability. Specifically, $d$ is discretized into 2m intervals up to 12m, $\ell$ into 50m intervals up to 300m, and $\beta$ into 5m intervals up to 30m, each with an open-ended bin beyond the upper bound. These cutoffs reflect common maritime conventions (Panama Canal Authority, 2022; Clarksons, 2025) (e.g., Panamax thresholds at $\ell\approx 300$m, $\beta\approx 32$m, and $d\approx 12$m) while keeping sufficient support.
@@ 4.2.2. Behavior Abstraction
This component converts motion signals within the minimal segment into a discrete, human-interpretable behavior pattern $v_{b}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$ and constrains the vocabulary growth for interpretability and efficient retrieval.
Kinematic Tokens $(p^{s},p^{\theta},p^{\psi})$. Speed $s$, course over ground $\theta$, and heading $\psi$ exhibit diverse dynamics; direct tokenization yields opaque categories and uncontrolled growth. An LLM-assisted procedure analyzes the temporal evolution in $\mathbf{S}_{\iota}^{k}$ and maps them to discrete variables $(p^{s},p^{\theta},p^{\psi})$. The vocabulary size is controlled by finite sets $\mathcal{P}^{s}$, $\mathcal{P}^{\theta}$, and $\mathcal{P}^{\psi}$ with a deduplication rule that reuses semantically equivalent tokens before adding new ones.
@@
Duration Token ($p^{\tau}$). Duration token $p^{\tau}$ is computed from the segment time span and discretized into fixed-width 50s intervals with an open-ended bin for long gaps (i.e., $3600,+\infty)$).
@@ 4.2.3. Imputation Method Builder
The Imputation Method Builder produces the imputation method $v_{f}=(f,d(f))$, where $f$ is an executable function and $d(f)$ is a concise usage description. Given a minimal segment $\mathbf{S}_{\iota}^{k}$, $v_{f}$ is generated through a three-step process.
Function Proposal. Based on $\mathbf{S}_{\iota}^{k}$ and the associated $(v_{s},v_{b})$, we first retrieve the most relevant imputation method $v_{f}\in\mathcal{V}_{f}$ using the retrieval scheme described in Section 4.3.2, to prevent unnecessary duplication and keep $\mathcal{V}_{f}$ compact. If no suitable candidate is found, an LLM-based function generator produces a new executable function $f$ that directly fits the observed trajectory.
Execution and Validation. The proposed function $f$ is executed on $\mathbf{S}_{\iota}^{k}$. At each timestamp we compute the absolute errors in latitude and longitude, then average them to obtain the overall fitting error $e(f)=\tfrac{1}{2}(\mathrm{MAE}_{\phi}+\mathrm{MAE}_{\lambda})$. The function is accepted if $e(f)$ does not exceed a predefined threshold of $\varrho_{f}=3e-3$. If the error exceeds this threshold, the LLM-based function generator is provided with diagnostic feedback (error summaries) and asked to refine its proposal. This is repeated at most $\varrho_{r}$ times (default 3).
@@ 4.2.4. SD-KG Integration and Maintenance
Node Integration. Since static attributes, behavior patterns, and imputation methods have already undergone de-redundancy during knowledge unit extraction, the integration process is straightforward. For each static attribute, if a newly generated token does not yet exist, a corresponding node is created and added to $\mathcal{V}_{s}$. Similarly, if the behavior pattern $v_{b}$ is not present in $\mathcal{V}_{b}$, or the imputation method $v_{f}$ is not present in $\mathcal{V}_{f}$, a new node is inserted.
Edge Maintenance. For each knowledge unit $u=(v_{s},v_{b},v_{f})$, edges $(v_{s},v_{b})$ and $(v_{b},v_{f})$ are updated. If the edge does not exist, it is created with an initial weight of $1$. If it already exists, the weight is incremented by $1$, capturing the frequency of observed co-occurrences. This ensures that the SD-KG preserves both structural completeness and statistically grounded associations while remaining compact and retrievable.
@@ 4.3. Knowledge-Driven Trajectory Imputation
With the SD-KG $\mathcal{G}_{d}$ in place, we formalize knowledge-driven trajectory imputation for an incomplete minimal segment $\mathbf{S}_{\iota}^{k}$. The procedure comprises three components: Behavior Estimator, which uses contextual constraints to query $\mathcal{G}_{d}$, rank candidates, and select the behavior pattern with supporting evidence; Method Selector, which retrieves an imputation method aligned with the chosen pattern, and executes it to reconstruct the gap; and Explanation Composer, which derives a human-friendly explanation by combining a regulatory rule cue with an operational protocol rationale from the induced subgraph of SD-KG and contextual information. In what follows, we detail each component.
@@ 4.3.1. Behavior Estimator
The Behavior Estimator operates in two steps: Candidate Set Generation, which queries the SD-KG to produce a ranked shortlist of plausible behavior patterns, and LLM-driven Selection and Rationale, which selects the final pattern and articulates a supporting rationale.
Candidate Set Generation. Given an incomplete segment $\mathbf{S}_{\iota}^{k}$ with context $(u_{\iota}^{-},u_{\iota}^{+})$, it infers the vessel-specific static attribute set $\mathcal{V}^{k}_{s}(\iota)$ from $(u_{\iota}^{-},u_{\iota}^{+})$, since static attributes vary slowly over time. It then queries the SD-KG to retrieve behavior candidates via static–behavior edges $C_{b}$:
@@
and it assigns each candidate a graph prior based on normalized log-additive support:
(9) $\pi(v_{b})\;=\;\frac{\displaystyle\sum_{v\in\mathcal{V}^{k}_{s}(\iota)}\!\log\!\big(w_{sb}(v,v_{b})+1\big)}{\displaystyle\sum_{v_{b}^{\prime}\in\mathcal{C}_{b}}\;\sum_{v\in\mathcal{V}^{k}_{s}(\iota)}\!\log\!\big(w_{sb}(v,v_{b}^{\prime})+1\big)},$
where $w_{sb}$ is the weight of edge $(v,v_{b},w)$. It keeps the top-$K$ candidates by $\pi(v_{b})$ as the shortlist $\mathcal{C}_{b}^{K}$ (default K=5K{=}5).
LLM-driven Selection and Rationale. It first constructs the induced subgraph $\mathcal{G}_{\iota,b}^{K}=\mathcal{G}_{d}[\mathcal{V}^{k}_{s}(\iota)\cup\mathcal{C}_{b}^{K}]$, serializes it in a graph description language (i.e., DOT (Wikipedia contributors, 2025a)), and supplies it together with the contextual behavior patterns $v_{b}^{-}$ and $v_{b}^{+}$ extracted from $u_{\iota}^{-}$ and $u_{\iota}^{+}$. It then selects a final pattern $v_{b}^{*}\in\mathcal{C}_{b}^{K}$ based on the LLM’s analysis of this information.
Alongside the selection, it also obtains a two-part rationale: i) graph support, which highlights the most informative edges in $\mathcal{G}_{\iota,b}^{K}$ and their weights that favor the choice; and ii) contextual justification, which explains the consistency with $v_{b}^{-}$, $v_{b}^{+}$, and the gap’s boundary conditions. These two parts constitute $\mathcal{J}_{\iota}^{k,b}$.